OpenStudy (anonymous):

Solve 2(x - 4) - 6(2x - 3) = 0 A. x = -4 B. x = -1 C. x = 0 D. x = 1

OpenStudy (shaik0124):

2x-8-12x+18=0 -10x+10=0 x=1